OpenStudy (anonymous):

A tower in Maintown sends radio signal a certain distance (in miles) according to the equation: x2 + y2 = 8,100. The Smith family lives 50 miles south and 70 miles east of Maintown. Part 1: Can the Smith family receive the radio signal from Maintown? (3 points) Part 2: In complete sentences, explain your reasoning. (3 points)

OpenStudy (mathstudent55):

Plug in 50 for x and 70 for y, and evaluate x^2 + y^2

OpenStudy (anonymous):

hmm...ok i plugged it in. but what do you mean by "evaluate"?

OpenStudy (mathstudent55):

Calculate a number

OpenStudy (mathstudent55):

x^2 + y^2 for x = 50 and y = 70: 50^2 + 70^2 =

OpenStudy (anonymous):

sorry for the long wait..is it 120^2?

OpenStudy (mathstudent55):

No. You need to square each number and add the squares together.

OpenStudy (mathstudent55):

50^2 + 70^2 = 2500 + 4900 = 7400

OpenStudy (anonymous):

oh...ok that's easy...so that's the answer?

OpenStudy (mathstudent55):

Since 7400 is less than 8100, they are within the range to get reception.

OpenStudy (anonymous):

thanks lot man :) peace bro.